the decibel level of sound is given by D= 10 log (I/10^-12), WHERE I is the sound intensity measured in watts per square meter. Find the decibel level of a
whisper at an intensity of 5.4 x 10-^10 watts per square meter.
so does that mean:
D= 10 log (5.4 x 10^-10)/ 10^-12)
do you subtract the exponents to get
so pluggin 5.4 x 10^-10 in a calculator i got 6.798
and 10^-12 I got 1.318
divide the two 6.798/1.318= 5.158
d=10 log (5.158) except that isn't one of my choices so i know i am wrong somewhere.
a. 2.73 decibels
b. 3.73 decibels
c. 27.3 decibels
d. 37.3 decibels




Answer by Nate(3500)

D = 10*log(I10^12)
D = 10*log((5.4)(10^-10)(10^12))
D = 10*log(540)
About 27.3239
c.